# Revolutionizing Large Language Model Training through Dynamic Parameter Adjustment

### v1

@@ 2 Methodology
A detailed analysis of the above results is postponed to Section 3.
3 Theoretical analysis
@@
In this section, we conduct a thorough discussion of our algorithm and address four key aspects:
@@
$\mathbf{y}=(\mathbf{W}+\frac{1}{r}\mathbf{B}\mathbf{A})\mathbf{x}.$ (5)
Next, we calculate the gradients of the column vectors of $\mathbf{B}$. Recall the decomposition of $\mathbf{B}\mathbf{A}$ as defined in the previous equations. For $k=1,\ldots,r$, the gradient of $\mathbf{b}_{k}$ with respect to the loss function $\mathcal{L}$ is given by:
@@
In our algorithm, we randomly select candidate vectors to replace vectors in $\mathbf{A}$ and $\mathbf{B}$, which alters the matching pairs of $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$. A natural question arises: Does the matching order of these vector pairs influence the training effects?
@@
To integrate the LoRA matrices while preserving the initial loss value, we reformulate $\mathcal{L}(\tilde{\mathbf{W}}\mathbf{x})$ as $\mathcal{L}((\mathbf{W}-\sum_{k}{\mathbf{b}}_{k}{\mathbf{a}}_{k}^{T}+\sum_{k}\tilde{\mathbf{b}}_{k}\tilde{\mathbf{a}}_{k}^{T})\mathbf{x})$. Further, we simplify this expression to $\mathcal{L}(\mathbf{a}_{1},\cdots,\mathbf{a}_{r};\mathbf{b}_{1},\cdots,\mathbf{b}_{k};\mathbf{x})$. A simple observation is
$\mathcal{L}(\mathbf{a}_{1},\cdots,\mathbf{a}_{r};\mathbf{b}_{1},\cdots,\mathbf{b}_{k};\mathbf{x})=\mathcal{L}(\mathbf{0},\cdots,\mathbf{0};\mathbf{0},\cdots,\mathbf{0};\mathbf{x}).$ (8)
@@
where $\Delta\mathbf{b}_{k}\Delta\mathbf{a}_{k}^{T}$ represents a minor term that can generally be disregarded. Hence, the updates derived by $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ are independent.
From this discussion, we can conclude that the order of vectors $\mathbf{a}_{k}$ and $\mathbf{b}_{k}$ does not influence the parameter updates in the current step. For instance, for $1\leq i,j\leq r$, back propagation of $\mathcal{L}(\mathbf{a}_{1},\cdots,\mathbf{a}_{j},\cdots,\mathbf{a}_{i},\cdots,\mathbf{a}_{r};\mathbf{b}_{1},\cdots,\mathbf{b}_{k};\mathbf{x})$ and $\mathcal{L}(\mathbf{a}_{1},\cdots,\mathbf{a}_{i},\cdots,\mathbf{a}_{j},\cdots,\mathbf{a}_{r};\mathbf{b}_{1},\cdots,\mathbf{b}_{k};\mathbf{x})$ yield the same parameters updating to the weight matrix of the linear layer.
Effectiveness of SwiLoRA
@@
We now compare the modified model with another model that implements the SwiLoRA strategy. Define $\mathbf{B}^{0}_{:,i}=\mathcal{C}(\mathbf{B})[i]$ and $\mathbf{A}^{0}_{i,:}=\mathcal{C}(\mathbf{A}^{T})[i]^{T}$ for $i=1,\ldots,\min(m,n)$. It becomes apparent that the two models are quite the same except that the model applying SwiLoRA strategy updates only subsets of parameters incrementally.
In optimization, it is well-established that for problems with separable objective functions, the parameters of each separable group can be optimized independently. Although the loss function of the SwiLoRA model is not separable, the preceding discussion has demonstrated the independence between the LoRA vectors. Consequently, we can infer that the inseparable components of the loss function concerning parameters within the same linear layer are modest. Therefore, this suggests that training subsets of parameters incrementally, as in the SwiLoRA model, is likely more effective than other methods, such as the layer-wise training approach proposed in [2].
@@
Derivation of initialization Parameters
We have provided the initial values of $\mathbf{B}$ and $\mathbf{A}$ in Section 2. Below, we present the deduction process.
@@
As demonstrated in [12], consider two matrices, $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, both characterized by zero mean and uniform distribution. The standard deviation (std) of the elements of their product is given by:
@@
Assuming the updated parameters are solely influenced by the gradients of the current step, to obtain $\frac{1}{r}\Delta BA\sim\frac{1}{r}B\Delta A$, the following condition must be met:
@@
By combining (13)-(3), we achieve the following standard deviations:

### v2

@@ 2.2 SwitchLoRA
A detailed analysis of the above results can be found in Appendix A.
3 Related work
Direct low-rank factorization method
Numerous studies [6, 44, 53, 23] have demonstrated the effectiveness of using low-rank factorization to approximate the weights of linear layers in deep neural networks. They employ methods such as SVD to achieve a factorization $\mathbf{U}\mathbf{V}$ that minimizes $\|\mathbf{W}-\mathbf{U}\mathbf{V}\|$. Later on, Pufferfish [49] and subsequent work in Cuttlefish [50] employ full-rank training prior to low-rank training to enhance efficiency. Additionally, they introduce adaptive strategies to determine the necessary duration of full-rank training and to select the appropriate rank for each linear layer for SVD. Further developments in this field include InRank [59], which proposes a low-rank training approach based on greedy low-rank learning [32]. Additional research such as [43] integrates orthogonality into the low-rank models to enhance training accuracy, while [19] introduces low-rank ordered decomposition, a generalization of SVD aimed at improving low-rank training efficiency.
These innovations mainly focus on convolutional neural networks (CNNs) and smaller-scale language models.
LoRA variants
After the introduction of LoRA in [21], which facilitated fine-tuning with very few trainable parameters, numerous works are proposed to improve the performance of LoRA. Improvements include better initialization strategies for LoRA matrices as demonstrated in [51, 52, 38]. Additionally, [16, 25] have adjusted learning rates for $\mathbf{B}$ and $\mathbf{A}$ to optimize training outcomes. Other research efforts, such as those in [27, 35], have modified the training process of LoRA. Moreover, some studies, such as [14, 58], focus on training models from scratch within a sparse model structure.
Similar to our approach, various LoRA variants employ strategies to increase the rank of updated parameters by merging parameters of adapters into $\mathbf{W}$. For instance, Chain of LoRA [54] and ReLoRA [34] merge $\mathbf{B}\mathbf{A}$ into $\mathbf{W}$ and restart training at regular intervals. ReLoRA enables low-rank training during the early phases, yet it still requires 33% of the steps to be full-rank training. Delta-LoRA [62], another variant, targets the fine-tuning phase by updating the matrix $\mathbf{W}$ using the gradients from the LoRA matrices $\mathbf{B}$ and $\mathbf{A}$ as they are updated, enhancing accuracy for fine-tuning.
Other compression methods
In addition to previously discussed techniques, there are many other methods to compress models during training. For instance, several studies have introduced quantization to LoRA [8, 31, 24], effectively reducing memory overhead during fine-tuning. Other research employs iterative pruning and growth techniques during training [11, 55, 37, 10]. Additionally, some works focus on compressing gradients through quantization [7, 28] or gradient projection [60]. Notably, [60] presents a recent method for training from scratch that utilizes SVD to project gradients into a periodically updated subspace. This approach also enables the addition of quantization, offering enhanced memory efficiency compared to LoRA.
4 Experiments
4.1 Experimental setup
Our studies are carried out on the LLaMA model [45], with model sizes reduced to 130M, 250M, and 350M. We designed our experiments based on the settings described in [34] to benefit from established hyperparameter configurations. The specific hyperparameters for these models are detailed in Table 1. We use Adam optimizer to train the model with $\beta_{1}=0.9,\beta_{2}=0.999$. We use a cosine learning rate schedule with 100 warm-up steps and a total of 40,000 training steps.
The pre-training experiments utilize the C4 dataset [9], with the first 46M samples of the training dataset serving as our training data, and samples from the entire validation dataset used for testing. The evaluation of validation loss is performed on 10M tokens for all our experiments, with evaluations conducted every 1,000 steps. Additionally, we utilize some of tasks from the GLUE benchmark [48] to assess the reasoning capabilities of the models. All experiments are conducted using 8xNVIDIA Tesla A100 80GB PCIe GPUs. Gradient accumulation is applied when GPU memory reaches its limit.
We have conducted ablation studies to assess the impact of various configurations, detailed in Appendix B.
Table 1: Model sizes and architectures used in our experiments
Params Hidden Heads Layers Batch size Batch size per GPU Seq. len.
130M 768 12 12 600 150 256
250M 768 16 24 1152 72 512
350M 1024 16 24 1152 72 512
1.3B 2048 32 24 1536 16 512
To ensure fairness across all experiments, the initialization method described in Section 2 is applied to both LoRA and SwitchLoRA experiments. We deploy LoRA adapters across all attention layers and fully connected layers in these experiments.
All experiments were repeated multiple times to select the best results. The learning rates for pre-training experiments were selected from a predefined set $\cup_{n=2,3,4}${1e-n 2e-n, 5e-n}. We have determined that the optimal learning rate remains consistent across different model sizes for all methods. Specifically, the learning rate for full-rank training is set at 0.001, while the learning rate for the LoRA method is 0.01. For SwitchLoRA, the learning rate is slightly higher at 0.02.
4.2 Basic experiments
Figures 2 displays the experimental results for the 130M, 250M, and 350M models, respectively, with the LoRA rank set to $128$. The data reveal that while LoRA alone does not yield satisfactory training results, SwitchLoRA approaches the performance of full-rank training. The performance gap continues to grow as model size increases. This suggests that the low-rank training approach, such as LoRA, might cause models to become trapped in local minima, while SwitchLoRA mitigates this issue by dynamically changing trainable parameters.
Table 2: Perplexity results at step 38,000 for 130M, 250M and 350M.
130M 250M 350M
Full-rank 27.71 20.19 18.72
LoRA(rank$=128$) 34.74 29.56 31.87
SwitchLoRA(rank$=128$) 30.26 20.97 19.96
[…]
Table 3: Perplexity results at step 38,000 for 1.3B models.
1.3B
Full-rank 15.23
SwitchLoRA(rank$=256$) 15.89
SwitchLoRA(rank$=512$) 15.01
Figure 3: Loss results for 250M, 350M and 1.3B models using higher LoRA ranks.
As shown in Figure 3, additional experiments conducted on the 250M, 350M and 1.3B models using higher LoRA ranks demonstrates improved performance compared to those with the rank set at $128$, achieving outcomes close to those of full-rank training. Although utilizing a higher rank yields better outcomes, it may not be more economical to increase the LoRA rank instead of increasing the model size for larger models for several reasons. First, the method still has potential for further refinement. Second, a lower LoRA rank enables training on devices with limited memory capacities. Furthermore, in the context of 3D parallelism, inter-node communication is predominantly influenced by data parallelism, where communication overhead is proportional to trainable parameters. The trainable parameters for each model are detailed in Table 4. For further discussions on potential ways to enhance the SwitchLoRA strategy, refer to Section 5. Additionally, the impact of distributed training is detailed in Appendix F.
Table 4: Comparison of trainable parameters: full-rank models vs. LoRA and SwitchLoRA.
Full-rank 247.5M 247.5M 368.2M 368.2M 1339.5M 1339.5M
(Switch)LoRA $r=128$ $r=256$ $r=128$ $r=256$ $r=256$ $r=512$
98.9M 148.4M 125.6M 185.4M 370.7M 609.7M
4.3 Comparison with other methods
Among all related methods, the works which are most close to ours are ReLoRA [34] and GaLore [60]. We do comparison experiments on these two methods to further validate the effectiveness of our algorithm. The learning rates for all methods are tuned in $\cup_{n=2,3,4}${1e-n 2e-n, 5e-n}.
Comparison with ReLoRA
Figure 4: Comparison between ReLoRA and SwitchLoRA. In the figure, red circles denotes the steps at which the parameters of the LoRA adapter are reset.
Since ReLoRA requires full-rank pre-training as warm-up, we do full-rank pre-training on SwitchLoRA too to do a fair comparison. We train 250M LLaMA model specified in Section 4, with detailed settings available in Appendix C. In the Figure 4, we compare ReLoRA and SwitchLoRA with different full-rank pre-training steps. It shows that our method can still perform better when ReLoRA uses 5,000 steps full-rank pre-training and SwitchLoRA uses 200 steps full-rank pre-training. Furthermore, when both algorithms are subjected to the same 1,000 steps of full-rank pre-training, SwitchLoRA shows significant improvements on ReLoRA.
The frequency for resetting the LoRA adapters in ReLoRA is set to 1/5,000, significantly lower than the initial switching frequency of 1/40 in SwitchLoRA experiments. As illustrated in Figure 4, we observe a rapid decrease in loss at each resetting step in the ReLoRA experiments. In contrast, the loss reduction in SwitchLoRA experiments is steady and more rapid.
Comparison with GaLore
In the comparison experiments with GaLore, we strictly follow the setup in Galore [60]. Detailed setup can be found in Appendix C. For the 350M LLaMA model, GaLore achieves a perplexity of 20.29, whereas SwitchLoRA performs slightly better, with a perplexity of 19.58. In addition, we conducted additional experiments on the 350M model, changing only one hyperparameter to assess its impact. The perplexity results are shown in Table 5.
When further reducing the rank, as shown in Table 5, our method performs significantly better. This improvement may be because GaLore’s use of SVD focuses on the most significant directions, whereas SwitchLoRA covers all update directions, including less important ones that still require training.
Table 5: Perplexity comparison for GaLore and SwitchLoRA with different experimental setup.
Standard Model size=130M Rank=128 Rank=32 Seq. len. = 512
GaLore 20.29 26.17 22.52 34.09 19.03
SwitchLoRA 19.58 25.93 20.93 25.26 18.19
The gradient projection subspace update frequency in GaLore is set at 1/200, while the initial switching frequency for SwitchLoRA is 1/40. Additionally, since updates in GaLore are performed via SVD, the subspace changes are less frequent compared to approaches that randomly select a new subspace. Consequently, the subspace changes in GaLore are, in fact, less efficient.
4.4 Reasoning ability comparison
Current works on low-rank training for LLMs, such as [34, 60], primarily evaluate models based on perplexity and lack validation of reasoning abilities. To validate the reasoning abilities, we also conducted full fine-tuning using the resulting checkpoints from the aforementioned experiments. We fine-tuned the models on GLUE tasks [48]. For the checkpoints trained using SwitchLoRA, all LoRA adapters are merged into the original weights such that $\mathbf{W}\leftarrow\mathbf{W}+\mathbf{B}\mathbf{A}$ before the fine-tuning process. Detailed experiment settings are provided in Appendix C.
Table 6: GLUE benchmark of the full-rank, SwitchLoRA and GaLore pre-trained 350M models. The metric for STS-B is the Pearson correlation, while Matthew’s correlation coefficient is used for CoLA. Accuracy is reported for the other tasks.
CoLA STS-B MRPC RTE SST2
Full-rank pre-trained 42.95$\pm$5 87.26$\pm$0.2 79.16$\pm$1 59.86$\pm$1 90.88$\pm$0.5
SwitchLoRA pre-trained 23.13$\pm$15 87.71$\pm$0.5 76.86$\pm$2 56.24$\pm$5 90.83$\pm$0.3
GaLore pre-trained 40.23$\pm$2 86.14$\pm$0.5 72.70$\pm$4 54.66$\pm$4 89.35$\pm$0.5
We first perform full fine-tuning on the pre-trained 350M models. The full-rank pre-trained model is from Section 4.2. Similarly, the SwitchLoRA pre-trained model, with a LoRA rank of 256, is also from Section 4.2. The GaLore pre-trained model originates from newly conducted experiments, where the batch size, sequence length, and rank are the same as in the SwitchLoRA experiment. This GaLore pre-training experiment resulted in a perplexity of 21.61.
The full fine-tuning results for these three models are shown in Table 6. From the results, we observe that for the 350M models, except for the CoLA task, SwitchLoRA outperforms GaLore by [insert performance difference], and outperforms the full-rank model by [insert performance difference].
We also conduct fine-tuning experiments on the 1.3B LLaMA models, one pre-trained using full-rank and the other pre-trained using SwitchLoRA with a LoRA rank of 512. Both pre-trained models are from Section 4.2. As shown in Table 7, SwitchLoRA performs slightly worse in some tasks and better in others compared to the full-rank results. Overall, the average score of SwitchLoRA exceeds the full-rank results by approximately 1%.
Table 7: GLUE benchmark of the Full-rank pre-trained and SwitchLoRA pre-trained models. The metrics for all tasks except STS-B are accuracy. The metric for STS-B is Pearson correlation coefficient.
CoLA STS-B MRPC RTE SST2
Full-rank pre-trained 48.60$\pm$2 87.64$\pm$0.1 78.43$\pm$1 58.05$\pm$3 91.93$\pm$1
SwitchLoRA pre-trained 47.43$\pm$3 88.49$\pm$0.3 80.15$\pm$2 61.37$\pm$3 92.39$\pm$0.5
5 Limitations and future work
While our results are promising, there are several areas for future exploration. In our experiments, we have demonstrated that selecting a larger LoRA rank is necessary to achieve accuracy comparable to full-rank training. Additionally, finely tuning the switching frequency of the LoRA vectors presents significant challenges. To address these limitations, we propose the following directions for future work, as illustrated in Figure 5.
  • In our experiments, we simply used exponentially decreasing switching frequencies, which may not be the optimal approach. Guidelines should be developed to help set appropriate switching frequencies throughout the training process.
  • Going further, a more detailed idea is to examine each layer of the model to adjust the switching frequencies. For instance, LoRA-drop [61] evaluates whether the rank is sufficient using a norm of $\Delta\mathbf{W}\mathbf{x}$. This is rational because different types of layers, such as the $Q,K,V$ matrices in transformer layers, exhibit significantly varied behaviors.
  • In our work, we simply chose candidate vectors at random or sequentially. However, during training, all candidates are updated separately, leading to significant differences among them. The selection of these candidates may improve the training outcomes.
Figure 5: Future work roadmap.
6 Conclusions
In this work, we introduce SwitchLoRA, a novel training strategy designed for parameter-efficient pre-training. Our approach achieves comparable accuracy to full-rank training while reducing the trainable parameters to approximately 50% to 60% of those in traditional full-rank training. Moreover, the computational overhead and memory usage are nearly identical to those of LoRA when using the same number of trainable parameters. We further validate the reasoning abilities of models trained with SwitchLoRA using the GLUE benchmark. The results from the 1.3B model indicate that SwitchLoRA not only matches but also slightly outperforms full-rank training by about 1% in accuracy.
Appendix A Theoretical analysis
@@
In this section, we conduct a thorough discussion of our algorithm and address the following key aspects:
@@
$\mathbf{y}=(\mathbf{W}+\frac{1}{r}\mathbf{B}\mathbf{A})\mathbf{x},$ (4)
where the bias term is omitted for simplicity.
Next, we calculate the gradients of the column vectors of $\mathbf{B}$. For a function $f(\mathbf{x})$, we denote $\nabla_{\mathbf{x}}f$ as the partial derivative of $f$ with respect to $\mathbf{x}$. Recall the decomposition of $\mathbf{B}\mathbf{A}$ as defined in the previous equations. For $k=1,\ldots,r$, the gradient of $\mathbf{b}_{k}$ with respect to the loss function $\mathcal{L}$ is given by:
@@
In fact, the gradients expressed in (5) and (6) and be derived as follows:
Consider the expression for $\mathbf{y}_{i}$ given by $\mathbf{y}_{i}=\sum_{j,k}\mathbf{B}_{ij}\mathbf{A}_{jk}\mathbf{x}_{k}+\sum_{j}\mathbf{W}_{ij}\mathbf{x}_{j}$ for $i=1,\ldots,m$. The partial derivative of the loss function $\mathcal{L}$ with respect to $\mathbf{B}_{ij}$ is computed as
$\frac{\partial\mathcal{L}}{\partial\mathbf{B}_{ij}}=\sum_{k}\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{k}}\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{B}_{ij}}=\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{i}}\frac{\partial\mathbf{y}_{i}}{\partial\mathbf{B}_{ij}}=\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{i}}\sum_{k}\mathbf{A}_{jk}\mathbf{x}_{k},$ (7)
where we use the fact that $\frac{\partial\mathbf{y}_{k}}{\partial\mathbf{B}_{ij}}=0$ when $k\neq i$. This derivation confirms (5). Similarly, the derivative with respect to $\mathbf{A}_{jk}$ is
$\frac{\partial\mathcal{L}}{\partial\mathbf{A}_{jk}}=\sum_{i}\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{i}}\frac{\partial\mathbf{y}_{i}}{\partial\mathbf{A}_{jk}}=\sum_{i}\frac{\partial\mathcal{L}}{\partial\mathbf{y}_{i}}\mathbf{B}_{ij}\mathbf{x}_{k}.$
This calculation leads to (6).
@@
In our algorithm, candidate vectors are either randomly selected or chosen sequentially to replace vectors in $\mathbf{A}$ and $\mathbf{B}$, which alters the matching pairs of $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$. A natural question arises: Does the matching order of these vector pairs influence the training effects?
@@
To integrate the LoRA matrices while preserving the initial loss value, we reformulate $\mathcal{L}(\tilde{\mathbf{W}}\mathbf{x})$ as $\mathcal{L}((\mathbf{W}-\sum_{k}{\mathbf{b}}_{k}{\mathbf{a}}_{k}^{T}+\sum_{k}\tilde{\mathbf{b}}_{k}\tilde{\mathbf{a}}_{k}^{T})\mathbf{x})$. Further, we simplify this expression to $\mathcal{L}(\mathbf{a}_{1},\ldots,\mathbf{a}_{r};\mathbf{b}_{1},\ldots,\mathbf{b}_{k};\mathbf{x})$. A simple observation is
$\mathcal{L}(\mathbf{a}_{1},\ldots,\mathbf{a}_{r};\mathbf{b}_{1},\ldots,\mathbf{b}_{k};\mathbf{x})=\mathcal{L}(\mathbf{0},\ldots,\mathbf{0};\mathbf{0},\ldots,\mathbf{0};\mathbf{x}).$ (8)
@@
where $\Delta\mathbf{b}_{k}\Delta\mathbf{a}_{k}^{T}$ represents a minor term that can generally be disregarded. Hence, the updates derived by $\mathbf{b}_{k}$ and $\mathbf{a}_{k}$ are nearly independent.
From this discussion, we can conclude that the order of vectors $\mathbf{a}_{k}$ and $\mathbf{b}_{k}$ does not influence the parameter updates in the current step. For instance, for $1\leq i,j\leq r$, back propagation of $\mathcal{L}(\mathbf{a}_{1},\ldots,\mathbf{a}_{j},\ldots,\mathbf{a}_{i},\ldots,\mathbf{a}_{r};\mathbf{b}_{1},\ldots,\mathbf{b}_{k};\mathbf{x})$ and $\mathcal{L}(\mathbf{a}_{1},\ldots,\mathbf{a}_{i},\ldots,\mathbf{a}_{j},\ldots,\mathbf{a}_{r};\mathbf{b}_{1},\ldots,\mathbf{b}_{k};\mathbf{x})$ yield almost the same parameters updating to the weight matrix of the linear layer.
Effectiveness of SwitchLoRA
@@
We now compare the modified model with another model that implements the SwitchLoRA strategy. Define $\mathbf{B}^{0}_{:,i}=\mathcal{C}(\mathbf{B})[i]$ and $\mathbf{A}^{0}_{i,:}=\mathcal{C}(\mathbf{A}^{T})[i]^{T}$ for $i=1,\ldots,\min(m,n)$. It becomes apparent that the two models are quite the same except that the model applying SwitchLoRA strategy updates only subsets of parameters incrementally.
In optimization, it is well-established that for problems with separable objective functions, the parameters of each separable group can be optimized independently. Although the loss function of the SwitchLoRA model is not separable, the preceding discussion has demonstrated the independence between the LoRA vectors. Consequently, we can infer that the inseparable components of the loss function concerning parameters within the same linear layer are modest. Therefore, this suggests that training subsets of parameters incrementally, as in the SwitchLoRA model, is likely more effective than other methods, such as the layer-wise training approach [3, 2].
@@
Derivation of parameters initialization
@@
The main idea of [12] and [17] is to maintain a balance in the variance of the activation and gradients across layers during forward and backward propagation. In this study, we focus on balancing the variance of activations. Furthermore, we aim to ensure the updated parameters derived from $\mathbf{B}$ are of the same amount as those derived from $\mathbf{A}$:
$\Delta\mathbf{B}\mathbf{A}\sim\mathbf{B}\Delta\mathbf{A}.$ (12)
Consider two matrices, $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, both characterized by zero mean and uniform distribution. The standard deviation (std) of the elements of their product is given by:
@@
Assuming the updated parameters are solely influenced by the gradients of the current step, to obtain (12), the following condition must be met:
@@
By combining (14)-(A), we achieve the following standard deviations: